OpenStudy (lovelyanna):

What is the slope of a line parallel to the line with equation 2x – 5y = 9? A. -9/5 B. -5/2 C. -2/5 D. 2/5

OpenStudy (lovelyanna):

I'm still confused.. :/

OpenStudy (campbell_st):

oops should be add 5y to both sides of the equation...what do you get..?

OpenStudy (lovelyanna):

Hmm.. I'm still confused, is it A?

OpenStudy (campbell_st):

ok... I'll attempt to start to make y the subject, this will get the line into slope intercept form. 2x -5y = 9 +5y +5y add 5 y to both sides of the equation --------------- 2x = 9 + 5y -9 -9 subtract 9 from both sides of the equation ---------------- 2x - 9 = 5y or 5y = 2x - 9 to get the line into slope intercept form divide every term by 5.. then the slope in the number infront of x

OpenStudy (shamil98):

Also, parallel lines have equal slopes.

OpenStudy (campbell_st):

then as @shamil98 said, parallel lines have the same slope. So once you have the slope for your given line, the parallel line has the same slope

OpenStudy (lovelyanna):

Is it D..?

OpenStudy (campbell_st):

correct... well done

OpenStudy (lovelyanna):

Thanks!